what is the diameter in nautical miles of the coverage area of a spot antenna on a geosynchronous satellite that has a beamwidth of 1.5 degrees?
Answer #1

The diameter of the coverage area is:

D=\alpha r

r is the distance of the geosynchronous orbit (35876Km)

1Km = 0.539957 Nautical miles

D=\frac{2\pi }{360}1.5(35786)0.539957

\boldsymbol{D=505.872\: Nautical\: Miles}